OpenStudy (horrorfreakk):

I need someone who's great at algebra to help me on this question!! Part A: Solve -np - 70 < 40 for n. Show your work. Part B: 4w - 7k = 28 for k. Show your work.

OpenStudy (denonakavro):

-np - 70 < 40 -np < 110

OpenStudy (denonakavro):

Then divide both sides by p -n<110/p divide both sides by -1 n<-110/p

OpenStudy (denonakavro):

4w - 7k = 28 -7k=28-4w k= (28-4w)/-7

OpenStudy (horrorfreakk):

Thank you.